# On an estimate on Götzky’s domain

### v1

@@ 1.2. Götzky’s domain
A fundamental domain for $\Gamma_{K}$ is a measurable set $F\subset\mathbb{H}^{n}$ if $\mathbb{H}^{n}=\bigcup_{\gamma\in\Gamma_{K}}\gamma(F)$ holds and (apart from a possible exceptional set of measure zero) no two points of $F$ are on the same $\Gamma_{K}$-orbit. Such a domain is described for any Hilbert modular group in [4]. Another fundamental domain was constructed by Götzky in [2] for the quadratic field $K=\mathbb{Q}(\sqrt{5})$. Although this latter construction works for any Euclidean quadratic field, in general it is only proved to contain a fundamental domain (the proof of this latter statement for $\mathbb{Q}(\sqrt{5})$ in [2] works for any quadratic Euclidean field).
@@
It may be surprising at first sight that these numerical computations did not support actual proofs of these estimates, since $\mathcal{F}_{d}$ is a straightforward generalization of the standard fundamental domain for the modular group $SL(2,\mathbb{Z})$ in $\mathbb{H}$. But in the latter case there is no analogue of the subgroup generated by $U_{d}$, and more importantly, the fundamental domain for the subgroup generated by the parabolic motions, i.e. the analogue of the set $\mathcal{S}_{d}$ looks substantially simpler, namely it is a strip bounded by two hyperbolic lines. By contrast, the sets $\mathcal{S}_{s_{1},s_{2}}^{d}$ look different for various values of $s_{1}$ and $s_{2}$ making the whole picture and the computations much more complicated.
@@ 2. Estimates on Götzky’s domain
from above on the set $\mathcal{S}_{s_{1},s_{2}}^{d}\cap\mathcal{T}$ where $s_{1},s_{2}>0$. For this end we will estimate on the set
@@
The key step towards the proof of Theorem 2.1 is the proper choice the parameter $a$. Once this problem is handled accurately at critical places, the other cases become easily treatable by a computer.
@@ 2.2. Outline of the proof
holds. Since $b>0$ it is enough to obtain such numbers $\lambda$ and $\mu$ that the roots of the quadratic polynomial on the right hand side of (1) are real and the smaller root is negative while the other one is at least the required bound. That is, we require
@@
    We are going to choose the parameter $a\in(-1;1)$ so that the maximum $g(a,b,c)$ of the function $f_{y_{1},y_{2}}(x_{1},x_{2})$ on $P_{a}$ is taken at vertex.
@@ 2.3. Restriction of the parameter $c$
As a first step we show that it is enough to prove the statement if $c\in[\varepsilon_{d}^{-1};\varepsilon_{d}]$. To this end we consider the map
@@
for any $n\in\mathbb{Z}$. Note that $\left|z_{1}z_{2}\right|^{2}=\left|(T_{n}z)_{1}(T_{n}z)_{2}\right|^{2}$ holds and $T_{n}$ takes the set $P_{a,d}^{s_{1},s_{2}}$ to
@@
Let $z\in\mathcal{S}_{d}\cap\mathcal{T}$ an arbitrary point, then $c\in[\varepsilon_{d}^{2k-1};\varepsilon_{d}^{2k+1}]$ for some $k\in\mathbb{Z}$. There is a $\nu\in\mathcal{O}_{K}$ such that $(z_{1}+\nu,z_{2}+\nu^{\prime})\in T_{-k}P_{a,d}^{\varepsilon_{d}^{k}y_{1},\varepsilon_{d}^{-k}y_{2}}$, and if $N(z)=N(z_{1},z_{2})=\left|z_{1}z_{2}\right|^{2}$, then (since $z\in\mathcal{S}_{d}\cap\mathcal{T}$) we get
@@
From here we restrict ourselves to the case $d=5$ and the parameter $d$ will mostly be omitted from the notations (e.g. we always write $\varepsilon$ instead of $\varepsilon_{d}$). In this section we prove the theorem in the cases when $c$ is close enough to one of the endpoints of the interval $[\varepsilon^{-1};\varepsilon]$. Observe first that even though the points $(z_{1},z_{2})=(x_{1}+iy_{1},x_{2}+iy_{2})$ considered in the following are not necessarily contained in $\mathcal{S}_{5}\cap\mathcal{T}$, yet for any of them holds that there are points in $\mathcal{S}_{5}\cap\mathcal{T}$ with the same $y_{1}$ and $y_{2}$. This means that the statement of Lemma 2.2 holds for them, i.e. we always have the bound $b>0.54$.
@@
To obtain an analogous result for $h_{y_{1},y_{2}}(x)$ one must be careful by the choice of $a$.
@@
The parameter $a$ will be chosen as a function of $c$. We choose different functions on different subintervals of $[\varepsilon^{-1};\varepsilon]$, a constant function will do on the middle intervals, while we have to be more precautious by the endpoints where we use linear functions.
Let us set $a=p(c-\varepsilon)+\frac{1}{\sqrt{d_{K}}}$ on the interval $[\varepsilon-\delta;\varepsilon]$ for some $p>0$ and $\delta>0$, and similarly, we set $a=p^{\prime}(c-\varepsilon^{-1})-\frac{1}{\sqrt{d_{K}}}$ on the interval $c\in[\varepsilon^{-1};\varepsilon^{-1}+\eta]$ for some $p^{\prime}>0$ and $\eta>0$. While a detailed analysis will be made in the former case, we simply choose $p^{\prime}=1$ in the latter which makes the computations less tedious and fortunately works.
@@
This expression can be seen as a function of $c$ with a fixed parameter $b$, let us denote its value by $g_{1}(b,c)$. In the following we assume that $b<0.56$ (otherwise the claim of the theorem holds) and then this function is strictly increasing on some interval $[\varepsilon-\delta;\varepsilon]$:
@@
For a fixed $0.560<b<$ 0.56 the function $g_{2}(b,c)$ is strictly decreasing on the interval $[\varepsilon^{-1};0.68]$.
@@ 2.5. Estimates on the middle intervals
that is clearly a lower bound for $H_{a}(c)$ on the interval $(0.68;1.48)$. Since $\tilde{H}_{a}(c)$ is increasing, it is enough to check $\tilde{H}_{a}(c)\geq 0$ at the left endpoint of each subinterval. Similarly, to estimate $g_{1}$ or $g_{2}$ it will be sufficient to do this at the endpoints once their derivative has a constant sign on a subinterval. These derivatives are
@@
As a first example we consider an interval $[1;1+\delta)$ and set $a=0$. As $\tilde{H}_{0}(1)=2/3$, we get that $H_{0}(c)\geq 0$ holds if $c\in[1;1+\delta]$. Next we consider the derivative of $\Delta_{a,b,c}$:
$\Delta_{a,b,c}^{\prime}=\frac{\sqrt{5}b}{2}[(a-1)-(a+1)c^{-2}]$
@@
We will choose $a$ such that $a<(c_{0}^{2}-1)/(c_{0}^{2}+1)$ holds and the value $D(a,c)$ is non-positive. The value $a$ that we get this way will be denoted by $a_{1}$. Once $a_{1}$ is chosen, we increase $c_{0}$ as in the first step above as much as possible to get $c_{1}$ and obtain the proof of the theorem for $c\in[c_{0};c_{1})$. Continuing in the same way determine the values $a_{2}<a_{3}<\dots$ and $c_{2}<c_{3}<\dots$ until we have $c_{n}\geq 1.48$ for some $n\in\mathbb{N}$, in which case we stop. We summarize this algorithm in the following steps:
@@
so it is enough to show that the right hand side is non-negative at the right endpoint. In accordance with this we work with the function $g_{1}(a,b,c)$, its derivative is increasing for every $a\in(-1;1)$. We check that this derivative is negative at $1$ (at the right endpoint) and so we can estimate by $g_{1}(a,b,c_{-1})$ (by the value at the left endpoint). Hence for $z\in\mathcal{S}_{5}\cap\mathcal{T}$ we have
@@ 2.6. The case of equality
Since $d_{K}=5$, from the inequalities $b>0.54$ and $c+c^{-1}\geq 2$ it follows that the coefficient of $w$ above is positive, hence $\tilde{g}^{\prime}$ is strictly increasing on $\mathbb{R}$ and takes the value 00 only once. (Note that e.g. in the case $d=2$ one cannot argue this way once $c$ is close to $1$ since the value $d_{K}=8$ is quite large.) So $\tilde{g}$ has only one local extremum, and this must be a local minimum since $\lim_{w\to\pm\infty}\tilde{g}(w)=\infty$. This means that independently of the choice of $a$ the function $\tilde{g}$ and then also $g$ take their maximum on the intervals above at one of the endpoints. $\Box$
@@
for $c\in[1;\varepsilon]$, hence it is enough to show that the right hand side above is positive for any $c\in[1.48;\varepsilon]$. To avoid the work with complicated algebraic expressions we check this via (finitely many) substitutions. (This method will also be applied in the subsequent proofs.) Namely, we show that the polynomial $F_{1}(c)=A_{1}(c)+0.56B_{1}(c)+8.001$ has $5$ roots, and therefore if $x_{0}$ is the biggest root, then $F_{1}$ must be strictly increasing on the interval $[x_{0};\infty)$. We do all this by giving pairs $c_{1},c_{2}$ of real numbers such that $c_{1}<c_{2}$ and the sign of $F_{1}(c_{1})$ and $F_{1}(c_{2})$ is different. One checks easily (e.g. by a computer) that
@@
One checks that $\tilde{\varphi}(0)<0$ and $\tilde{\varphi}(\varepsilon^{-2})=-3+\sqrt{5}<0$, so $\tilde{\varphi}(t)$ is negative for any $t\in[0;\varepsilon^{-2}]$ and hence (4) is proved. As $b>0.5$ we have $f(t)\leq\sqrt{5}t(\sqrt{5}t-1)(\sqrt{5}t-2)(t+\varepsilon^{-1})+4\varphi(t)=:\tilde{f}(t)$. Since $\tilde{f}(0)=0$, it is enough to show that $\tilde{f}^{\prime}$ is negative on the interval $[0;\varepsilon^{-2}]$. A computation gives $\tilde{f}^{\prime}(t)=20\sqrt{5}t^{3}+3\sqrt{5}\varepsilon^{-2}t^{2}+(47-27\sqrt{5})t+9\sqrt{5}-21$. Now the inequalities

### v2

@@ 1.2. Götzky’s domain
A fundamental domain for $\Gamma_{K}$ is a measurable set $F\subset\mathbb{H}^{n}$ such that $\mathbb{H}^{n}=\bigcup_{\gamma\in\Gamma_{K}}\gamma(F)$ holds and (apart from a possible exceptional set of measure zero) no two points of $F$ are on the same $\Gamma_{K}$-orbit. Such a domain is described for any Hilbert modular group in [9]. Another fundamental domain was constructed by Götzky in [5] for the quadratic field $K=\mathbb{Q}(\sqrt{5})$. Although this latter construction works for any Euclidean quadratic field, in general it is only proved to contain a fundamental domain (the proof of this latter statement for $\mathbb{Q}(\sqrt{5})$ in [5] works for any quadratic Euclidean field).
The shape of such domains for quadratic fields have been studied by Cohn [1, 2], Deutsch [3, 4], Jespers, Kiefer and del Río [7], and Quinn and Verjovsky [8]. More recently, a general reduction algorithm was given by Strömberg [10] for Hilbert modular groups over arbitrary totally real number fields.
@@
The height plays the same role here as the imaginary part of a point on the complex upper-half plane when the action of the group $SL(2,\mathbb{R})$ is considered. In [9] it is used to divide the fundamental domain of a Hilbert modular group into disjoint parts (cusp regions). While in the Euclidean case there is only one such region, lower bounds on the height still have some significance in such cases, e.g. they are used in [6] for the computation of totally elliptic conjugacy classes (an element of $\Gamma_{K}$ is totally elliptic, if the trace of every coordinate of it has absolute value less than $2$), or they can affect implied constants in other estimates, see e.g. Lemma 1.2.5 and Lemma 2.4.1 and their proofs in the thesis [11] of the author. The aforementioned lemmata have far-reaching consequences in estimates of automorphic forms, and though the implied constants and lower bounds on the height are important rather from the computational point of view, it may be desirable to have precise results at least in such a classical example like Götzky’s domain.
It can be surprising at first sight that the numerical computations in [4] did not support actual proofs of the conjectured estimates, since $\mathcal{F}_{d}$ is a straightforward generalization of the standard fundamental domain for the modular group $SL(2,\mathbb{Z})$ in $\mathbb{H}$. But in the latter case there is no analogue of the subgroup generated by $U_{d}$, and more importantly, the fundamental domain for the subgroup generated by the parabolic motions, i.e. the analogue of the set $\mathcal{S}_{d}$ looks substantially simpler, namely it is a strip bounded by two hyperbolic lines. By contrast, the sets $\mathcal{S}_{s_{1},s_{2}}^{d}$ look different for various values of $s_{1}$ and $s_{2}$ making the whole picture and the computations much more complicated.
@@ 2. Estimates on Götzky’s domain
from above on the set $\mathcal{S}_{s_{1},s_{2}}^{d}\cap\mathcal{T}$ where $s_{1},s_{2}>0$. To this end we will estimate on the set
@@
The key step towards the proof of Theorem 2.1 is the proper choice of the parameter $a$. Once this problem is handled accurately at critical places, the other cases become easily treatable by a computer.
@@ 2.2. Outline of the proof
holds. Since $b>0$ it is enough to obtain numbers $\lambda$, $\mu$ such that the roots of the quadratic polynomial on the right hand side of (1) are real and the smaller root is negative while the other one is at least the required bound. That is, we require
@@
    We are going to choose the parameter $a\in(-1;1)$ so that the maximum $g(a,b,c)$ of the function $f_{y_{1},y_{2}}(x_{1},x_{2})$ on $P_{a}$ is taken at a vertex.
@@ 2.3. Restriction of the parameter $c$
As a first step we show that it is enough to prove the statement of Theorem 2.1 if $c\in[\varepsilon_{d}^{-1};\varepsilon_{d}]$. To this end we consider the map
@@
for any $n\in\mathbb{Z}$. (recall that $(\varepsilon_{d}^{n})^{\prime}$ denotes the conjugate of $\varepsilon_{d}^{n}$ in $\mathbb{Q}(\sqrt{d})$). Note that $\left|z_{1}z_{2}\right|^{2}=\left|(T_{n}z)_{1}(T_{n}z)_{2}\right|^{2}$ holds and $T_{n}$ takes the set $P_{a,d}^{s_{1},s_{2}}$ to
@@
Let $z\in\mathcal{S}_{d}\cap\mathcal{T}$ be an arbitrary point, then $c\in[\varepsilon_{d}^{2k-1};\varepsilon_{d}^{2k+1}]$ for some $k\in\mathbb{Z}$. There is a $\nu\in\mathcal{O}_{K}$ such that $(z_{1}+\nu,z_{2}+\nu^{\prime})\in T_{-k}P_{a,d}^{\varepsilon_{d}^{k}y_{1},\varepsilon_{d}^{-k}y_{2}}$, and if $N(z)=N(z_{1},z_{2})=\left|z_{1}z_{2}\right|^{2}$, then (since $z\in\mathcal{S}_{d}\cap\mathcal{T}$) we get
@@
From here we restrict ourselves to the case $d=5$ and the parameter $d$ will mostly be omitted from the notations (e.g. we always write $\varepsilon$ instead of $\varepsilon_{d}$). Note that we have $d_{K}=5$, $\varepsilon=\frac{1+\sqrt{5}}{2}$ and $\varepsilon^{-1}=-\varepsilon^{\prime}=\frac{\sqrt{5}-1}{2}$ in this case.
In this section we prove the theorem in the cases when $c$ is close enough to one of the endpoints of the interval $[\varepsilon^{-1};\varepsilon]$. Observe first that even though the points $(z_{1},z_{2})=(x_{1}+iy_{1},x_{2}+iy_{2})$ considered in the following are not necessarily contained in $\mathcal{S}_{5}\cap\mathcal{T}$, yet for any of them holds that there are points in $\mathcal{S}_{5}\cap\mathcal{T}$ with the same $y_{1}$ and $y_{2}$. This means that the statement of Lemma 2.2 holds for them, i.e. we always have the bound $b>0.54$.
@@
To obtain an analogous result for $h_{y_{1},y_{2}}(x)$ one must be careful with the choice of $a$.
@@
The parameter $a$ will be chosen as a function of $c$. We choose different functions on different subintervals of $[\varepsilon^{-1};\varepsilon]$, a constant function will do on the middle intervals, while we have to be more precautious at the endpoints where we use linear functions.
Let us set $a=p(c-\varepsilon)+\frac{1}{\sqrt{d_{K}}}$ on the interval $c\in[\varepsilon-\delta;\varepsilon]$ for some $p>0$ and $\delta>0$, and similarly, we set $a=p^{\prime}(c-\varepsilon^{-1})-\frac{1}{\sqrt{d_{K}}}$ on the interval $c\in[\varepsilon^{-1};\varepsilon^{-1}+\eta]$ for some $p^{\prime}>0$ and $\eta>0$. While a detailed analysis will be made in the former case, we simply choose $p^{\prime}=1$ in the latter which makes the computations less tedious and fortunately works.
@@
This expression can be seen as a function of $c$ with a fixed parameter $b$, let us denote its value by $g_{1}(b,c)$. In the following we assume that $b<0.56$ (otherwise the claim of Theorem 2.1 holds) and then this function is strictly increasing on some interval $[\varepsilon-\delta;\varepsilon]$:
@@
For a fixed $0.560<b<$ 0.56 the function $g_{2}(b,c)$ is strictly decreasing in $c$ on the interval $[\varepsilon^{-1};0.68]$.
@@ 2.5. Estimates on the middle intervals
that is clearly a lower bound for $H_{a}(c)$ on the interval $(0.68;1.48)$. Since $\tilde{H}_{a}(c)$ is increasing, it is enough to check $\tilde{H}_{a}(c)\geq 0$ at the left endpoint of each subinterval. Similarly, to estimate $g_{1}$ or $g_{2}$ it will be sufficient to do this at the endpoints once their derivatives viewed as functions in $c$ have constant signs on some subintervals. These derivatives are
@@
As a first example we consider an interval $[1;1+\delta)$ and set $a=0$. As $\tilde{H}_{0}(1)=2/3$, we get that $H_{0}(c)\geq 0$ holds if $c\in[1;1+\delta]$. Next we consider the derivative of $\Delta_{a,b,c}$ with respect to $c$:
$\frac{\partial\Delta_{a,b,c}}{\partial c}=\frac{\sqrt{5}b}{2}[(a-1)-(a+1)c^{-2}]$
@@
We will choose $a$ such that $a<(c_{0}^{2}-1)/(c_{0}^{2}+1)$ holds and the value $D(a,c_{0})$ is non-positive. The value $a$ that we get this way will be denoted by $a_{1}$. Once $a_{1}$ is chosen, we increase $c_{0}$ as in the first step above as much as possible to get $c_{1}$ and obtain the proof of the theorem for $c\in[c_{0};c_{1})$. Continuing in the same way determine the values $a_{2}<a_{3}<\dots$ and $c_{2}<c_{3}<\dots$ until we have $c_{n}\geq 1.48$ for some $n\in\mathbb{N}$, in which case we stop. We summarize this algorithm in the following steps:
@@
so it is enough to show that the right hand side is non-negative at the right endpoint. In accordance with this we work with the function $g_{1}(a,b,c)$, its derivative with respect to $c$ is increasing for every $a\in(-1;1)$. We check that this derivative is negative at $1$ (at the right endpoint) and so we can estimate by $g_{1}(a,b,c_{-1})$ (by the value at the left endpoint). Hence for $z\in\mathcal{S}_{5}\cap\mathcal{T}$ we have
@@ 2.6. The case of equality
Since $d_{K}=5$, from the inequalities $b>0.54$ and $c+c^{-1}\geq 2$ it follows that the coefficient of $w$ above is positive,
hence $\tilde{g}^{\prime}$ is strictly increasing on $\mathbb{R}$ and takes the value 00 only once. (Note that e.g. in the case $d=2$ one cannot argue this way once $c$ is close to $1$ since the value $d_{K}=8$ is quite large.) So $\tilde{g}$ has only one local extremum, and this must be a local minimum since $\lim_{w\to\pm\infty}\tilde{g}(w)=\infty$. This means that independently of the choice of $a$ the function $\tilde{g}$ and then also $g$ take their maximum on the intervals above at one of the endpoints. $\Box$
@@
for $c\in[1;\varepsilon]$, hence it is enough to show that the right hand side above is positive for any $c\in[1.48;\varepsilon]$. To avoid the work with complicated algebraic expressions we check this via (finitely many) substitutions. (This method will also be applied in the subsequent proofs.) Namely, we show that the polynomial $F_{1}(c)=A_{1}(c)+0.56B_{1}(c)+8.001$ has $5$ roots, and therefore if $x_{0}$ is the biggest root, then $F_{1}$ must be strictly increasing on the interval $[x_{0};\infty)$. We do all this by giving pairs $c_{1},c_{2}$ of real numbers such that $c_{1}<c_{2}$ and the signs of $F_{1}(c_{1})$ and $F_{1}(c_{2})$ are different. One checks easily (e.g. by a computer) that
@@
One checks that $\tilde{\varphi}(0)<0$ and $\tilde{\varphi}(\varepsilon^{-2})=-3+\sqrt{5}<0$, so $\tilde{\varphi}(t)$ is negative for any $t\in[0;\varepsilon^{-2}]$ and hence (4) is proved.
As $b>0.5$ we have $f(t)\leq\sqrt{5}t(\sqrt{5}t-1)(\sqrt{5}t-2)(t+\varepsilon^{-1})+4\varphi(t)=:\tilde{f}(t)$. Since $\tilde{f}(0)=0$, it is enough to show that $\tilde{f}^{\prime}$ is negative on the interval $[0;\varepsilon^{-2}]$. A computation gives $\tilde{f}^{\prime}(t)=20\sqrt{5}t^{3}+3\sqrt{5}\varepsilon^{-2}t^{2}+(47-27\sqrt{5})t+9\sqrt{5}-21$. Now the inequalities